OpenStudy (sonya):

INIK HOW DID YOU GET 2X=2 5/3

OpenStudy (saifoo.khan):

find the value of x first.

OpenStudy (anonymous):

3x-1=4 3x=4+1 x=5/3 multiply 2 to get 2x 2x=2(5/3) 2x=10/3